) A 2000 kg truck is going 20 m/s collides with a brick wall bringing it to a stop. The truck is totaled, but the wall is fairly undamaged. If the collision lasts 100ms, find the average force exerted on the truck by the brick wall. How does this compare to the force that the truck exerts on the wall?
Answer #1

let average force exerted on truck by wall is F_{av}

then F_{av}*t=Mu-Mv

where (Mu-Mv)  is the change in momentum of truck

u = 20 m/s , v = 0 m/s

F_{av}*100*10^{-3}=2000*20-2000*0=40000

F_{av}=4*10^{5}N

this force is equal and opposite to the force that truck exerts on wall